A dedicated sports-car enthusiast polishes the inside and outside surfaces of a hubcap that is a section of a sphere. When shPlease help
Answer #1

Using curved mirror equations u as object distance, v as image distance and f as focal length

for first case

1/v - 1/u = 1/f
1/25 - 1/u = 1/f

for second case

1/v + 1/u = 1/f
1/10 + 1/u = 1/f

equating both we get

1/10 - 1/u = 1/25 + 1/u

1/10 - 1/25 = 2/u

u = 33.33 cm

put in 1 to get f

f = 7.69 cm
but R = 2f = 15.38cm